Most staircase in use today have 9 inch treads and 8.5 inch risers. What is the slope of a staircase with these measurements?
Answer by rfadrogane(214)

Most staircase in use today have 9 inch treads and 8.5 inch risers. What is the slope of a staircase with these measurements?
Sol'n:
let: m - the slope of the staircase
m = riser/tread = rise/run = (y2-y1)/(x2-x1)
m = 8.5/9
m = 17/18 ----answer